# RealityCraft: An In-Situ CAD+CAM Interface for Novices via Scene-Aware Augmented Reality

### v1

@@ 2.3. Novice Support Tools and Systems
Designing systems for novices requires consideration of simplicity, intuitiveness, and ease of learning. In the context of digital design and fabrication, this often translates to reducing the complexity of traditional CAD tools and making them more accessible. Tinkercad, for instance, has been successful in this regard by providing an entry-level CAD tool that uses simple geometric shapes and a drag-and-drop interface that is easy for beginners to grasp (Stemasov et al., 2021). Other systems for design and fabrication using simplified core building blocks include FlatFitFab (McCrae et al., 2014), which connects planar parts for a 3D model, and similarly Kyub (Baudisch et al., 2019), which follows a simpler block-based design direction. CraftML (Yeh and Kim, 2018) uses declarative syntax where the code dictates the output 3D structure. Work by Lee et al. (Lee et al., 2018, 2016) uses pose and movement detection for personalized furniture design and work by Tian et al. simplifies analog fabrication with woodcutter (Tian et al., 2018) and lathe (Tian et al., 2019) tools. Additionally, Suzuki et al. (Suzuki et al., 2022) discuss how virtual twins are instrumental in visualizing simulated behaviors when integrated with spatial references. The work of Pfeuffer et al. (Pfeuffer et al., 2017) shows the use of virtual buttons on the wrists for ease of access to interactable menus.
Beyond simplifying interfaces, there is a growing trend toward educational integration, where tools are designed not only for practical use but also as learning aids. Augmented Math by Chulpongsatorn et al. and Augmented Physics by Gunturu et al. generate interactive simulations by extracting and animating static diagrams from math and physics textbooks with AR. Blocks-to-CAD (Lafreniere and Grossman, 2018) is developed as a cross-bridge application for teaching a more complex design tool via a simpler, more gamified tool. Similarly, research by Cao et al. (Cao et al., 2023) highlights the use of VR and AR in educational settings, where these technologies provide immersive learning experiences that help users understand complex concepts more easily. Similar systems utilizing these technologies lower the learning barrier while boasting creativity (Jeong et al., 2023).
Furthermore, retrieving existing designs (Stemasov et al., 2020; Liang et al., 2022) and remixing (Follmer and Ishii, 2012; Follmer et al., 2010; Stemasov et al., 2023b; Roumen et al., 2018) enhances accessibility while balancing expressivity (Stemasov et al., 2021). ShapeFindAR by Stemasov et al. enables searching for future physical artifacts through spatial queries and text (Stemasov et al., 2022).
@@
Generative design has also been instrumental in achieving desired 3D outcomes with minimal effort. Some tools allow users to design and fabricate 3D models based on 2D sketches (Saul et al., 2010; Li et al., 2020; Kazi et al., 2017; Johnson et al., 2012; Igarashi et al., 1999), thereby reducing the complexity associated with directly creating models in 3D space. Other ways of generative design utilize speech inputs (Ballagas et al., 2019) and constrained parametric design (Shugrina et al., 2015).
By examining these areas, our work in RealityCraft builds upon the existing knowledge and practices in XR, in-situ tools, and novice-friendly systems, aiming to integrate these elements into a cohesive framework that enhances not only the design and fabrication process but also the learning experience.
@@ 3. DESIGN RATIONALE
When developing RealityCraft, several key design considerations were prioritized to ensure the system was both effective and intuitive for novice users. The main considerations included are user-friendly design and interaction, workspace awareness, and seamless integration:
@@ 3.1. User-Friendly Design and Interaction
To enhance the user experience in RealityCraft, we prioritize both simplicity in design and intuitive interaction. Our approach combines straightforward design logic with natural user interactions to create an accessible and engaging environment for all users. The system’s design and interaction features are crafted to ensure that even novice users can seamlessly create and manipulate models, while advanced users benefit from an efficient design process.
@@
RealityCraft is designed with a focus on simplicity to ensure a novice-friendly user experience. By utilizing CSG operations, users can easily create complex models by combining basic shapes (e.g., cuboids and spheres) through simple boolean operations such as union and subtraction. This approach reduces complexity, allowing users to design without managing technical parameters. The system categorizes shapes into solid and hole types, enabling the addition or removal of material to create more advanced designs effortlessly.
@@
Interactivity in RealityCraft is driven by natural hand gestures, recognized by Quest’s tracking system. These gestures allow users to manipulate objects directly, making the design process intuitive and engaging. Building upon the methods outlined in prior works (Kim et al., 2005; Weichel et al., 2014; Stemasov et al., 2024), RealityCraft lets users interact with the elements directly with their hands and without the need for external tools other than the headset. For instance, a pinching motion grabs and moves objects, while wrist rotation enables feature toggling, offering a seamless and immersive experience. Additionally, wrist-mounted menus provide an easy-to-use interface and buttons for quick navigation, visually and audibly confirming actions, enhancing usability.
@@ 3.2. Workspace Awareness
RealityCraft’s real-world integration is achieved through sophisticated scene recognition and depth occlusion technologies. These technologies allow the system to not only understand the environment but also intelligently interact with it.
Four photos labeled (a), (b), and two in (c). The first two photos show the first-person view of the user looking at a room with a table inside. The other two photos show three small decorative cacti on a table with a digital pyramid model behind them. In image (a), three yellow rays come out from the user’s right hand, first hitting the floor where a text appearing above says ”select: floor”, second hitting the top of the table where a text appearing above says ”select: table”, and the third hitting the wall where a text appearing above says ”select: wall”. There is a white outline around the table. In image (b) the room is seen from the same angle as image (a), but this time there are grids on top of the floor, table, and wall. In the first image of (c), the digital pyramid model partially blocks the view of the cacti even though it is placed behind them while in the second image of (c) the cacti are fully visible.
@@
It is essential for the system to recognize and adapt to its physical surroundings, using real-time environment mapping to adjust designs according to available space and nearby objects. The system can identify furniture and other objects, categorizing them as potential design spaces or obstacles such as a desk, wall, the ground, or similar surfaces. This recognition helps in placing and scaling models accurately within the environment. In our case, this reflects the system being aware of the selected workspace, a grid layout placed on top of the selected workspace helps users utilize snap-to-grid features and see the projection of the objects on the workspace, providing a reference while designing.
@@
It is also essential for users to understand their surroundings for context awareness while designing. While AR functionalities significantly aid in this area, models rendered in front of every world object can detract from the realistic nature we aim to achieve in the design task. By understanding the depth of objects in the environment, RealityCraft can occlude parts of a model that are behind physical objects, enhancing the realism and immersion of the design experience while elevating the user’s perception of the design as part of their real environment.
@@ 3.3. Seamless Integration
The transition between designing in AR and preparing a model for fabrication (in our case 3D printing) should be fluid and straightforward, encouraging users to iterate on designs quickly and efficiently.
@@
In RealityCraft, we have consolidated design and fabrication functionalities into a single system, streamlining the entire workflow. Users can design their models and select fabrication options directly within the AR environment, eliminating the need to transfer designs between separate tools. This integration simplifies the process by allowing users to handle all aspects of design and preparation within a unified interface, thereby reducing the complexity traditionally associated with using multiple software packages.
@@ 4. Implementation
RealityCraft was implemented for Meta Quest 3. The system code is written in C# with Unity Version 2022.3.22f1 LTS666https://unity.com/releases/2022-lts, Accessed: 19.07.2024. The system utilizes open-source resources provided by Meta to support various features. Specifically, we used Meta XR All-in-One SDK777https://developer.oculus.com/downloads/package/meta-xr-sdk-all-in-one-upm/, Accessed: 19.07.2024, which includes libraries for hand tracking, room and furniture detection (Scene API888https://developer.oculus.com/documentation/native/android/mobile-scene-api-ref/, Accessed: 2.08.2024), passthrough features, and other essential XR development tools for Unity. Additionally, we incorporated Meta’s Depth API999https://github.com/oculus-samples/Unity-DepthAPI, Accessed: 19.07.2024 library for the depth occlusion feature and Unity’s built-in audio engine for the spatial audio utilities. The subsequent sections will outline the envisioned procedural steps of using RealityCraft with its features, from (1) selecting the workspace to (2) designing artifacts, and finally to (3) fabrication of those artifacts.
@@ 4.1. Workspace Selection and Modification
RealityCraft system enables users to select their desired workspace, which can vary from the top of a desk to the floor, walls, door frames, and other flat surfaces.
Five photos labeled (a), (b), (c), (d), (e) show the first-person view of a user in front of a wall interacting with their hand and the digital objects. In images (a), (b), and (c) user have extended their left hand to their eye level and looking at the inside of their wrist. There are 12 white square buttons on the wrist in a three-by-four grid. Seven of the buttons show primitive geometric objects, namely cube, cylinder, sphere, capsule, triangular prism, pyramid, and cone. In the image (a), the user is hovering his right index finger on their wrist. In the image (b), the user pokes the cube button with their right index finger and the button turns green and a cube appears above the left hand. In the image (c), the user does the same with a sphere next to the cube. In image (d), user does a pinch motion while their right hand is directed toward the created cube and their left hand is oriented palm-down. The cube’s color is changed from blue to cyan and a white grid is appearing on the cube. Similarly, in the image (e), the user does the same pinch motion for the sphere, this time with the left hand oriented palm-up. Both of the objects turn to cyan, with white grid appearing on the sphere instead of the cube.
Six photos labeled (a), (b), (c) each label corresponding to two photos show a cube being manipulated with their right hand from the first-person view of the user in front of a wall. The first pair of pictures labeled (a) shows the user pinching the movement handle on top of the cube in the first picture and dragging their right hand and the object to the right in the second picture. The second pair of pictures labeled (b) shows the user pinching the rotation handle next to the cube in the first picture and rotating their right wrist and the cube in the second picture. The third pair of pictures labeled (c) shows the user pinching the corner handle at the top right corner of the cube in the first picture and moving their hand to the right and the cube stretching to the right in the second picture.
@@ 4.1.2. Workspace Selection
After scanning, users can interact with the RealityCraft system and view the highlighted selectable workspaces. This is possible with the headset’s built-in room recognition feature and MRUK framework101010https://developer.oculus.com/documentation/unity/unity-mr-utility-kit-overview/, Accessed: 20.07.2024. Then, users use their right hand to direct a ray toward the surfaces of furniture or room boundaries, which displays the name of the planned workspace it intersects (Figure 2a). An index finger pinch gesture is then employed to confirm the selection.
@@ 4.2.1. Object Creation
Users can choose from seven types of primitive 3D geometries (cube, sphere, cylinder, capsule, triangular prism, pyramid, and cone) using buttons on their left wrist (Figure 3a). The color of the button turns blue when the finger hovers close to the button. This gives the user feedback on which button is about to be pressed. Upon pressing a button, the selected object appears on the user’s left hand (Figure 3b and 3c).
@@ 4.2.2. Object Highlighting and Selection
While interacting with the system, users may encounter multiple instances of objects, necessitating an efficient method for selecting and manipulating specific items. The following mechanisms are then utilized.
@@
Three photos labeled (a), (b), and (c) show a first-person view of a user looking at a wall and interacting with a complex 3D abstract design. In images (a) and (b), the user has their left hand’s palm looking right and they apply a pinching motion to the corner handle on the upper right side of the manipulation box. A pink ray is seen coming from the middle and going through the selected corner handle. The user brings their right hand in the second picture closer to the center of the manipulation box, resulting in a smaller object with the same aspect ratio. Image (c) shows the user dragging the same corner handle with their left-hand palm looking down. The resulting object is compressed vertically.
@@
If the user makes a pinch motion (thumb and index finger touching) while an object is highlighted, the object will be selected (Figure 3d). Once selected, the object will change to a bright cyan color, and a sound will emanate from the direction of the selected object. An Object Manipulation Box will appear around the selected objects. The Object Manipulation Box features interactable elements for moving, resizing, and rotating the objects. It is essentially a bounding rectangular prism, maintaining a fixed rotation aligned with the default world axis and perfectly encapsulating the selected objects.
@@ 4.2.3. Single and Multiple Selection
Users can switch between two modes of object selection: single selection and multiple selection. These modes can be toggled based on the orientation of the left hand. If the left hand’s palm is facing down, the mode is single selection (Figure 3d); if the palm is facing up, the mode is multiple selection (Figure 3e).
@@
The single selection mode is straightforward. In this mode, only one object can be selected at a time, meaning that all the other selected objects will be deselected. As single selection mode is considered the default option, the palm-down orientation is chosen as it is the resting position of the hand.
@@
Multiple selection mode is similar to pressing the control key on a keyboard while selecting folders in a file browser. In this mode, selecting an object that is not already selected adds it to the list of selected objects. Conversely, selecting an already selected object removes it from the list. When multiple objects are selected they can be manipulated in synchronization.
@@ 4.2.4. Object Manipulation
With the help of the elements within the Object Manipulation Box appearing around the selected objects, users can move the objects freely, resize them in all directions, and rotate them along different axes.
@@
Moving objects is done with the help of a movement handle. This handle appears on top of the selected objects, and when picked up with a pinch gesture, it follows the tip of the fingers on the pinch location (Figure 4a). This enables completely free movement, and users can move objects around with the exact movement of their hands.
@@
Rotating objects involves rotating them around the primary axes (x, y, z) of the Object Manipulation Box. These axes extend perpendicularly through the sides of the box from its center. Along these axes, there are placed three rotation handles, each responsible for rotation around one specific axis. When a rotation handle is grasped with a pinch gesture, it functions similarly to a knob, allowing free rotation along the corresponding axis based on the hand’s movement (Figure 4b).
@@
The size of the selected objects can be adjusted using the corner handles located at each vertex of the Object Manipulation Box (Figure 4c). When a user grabs a corner handle with a pinch gesture, it changes color and shape to provide visual feedback, indicating that the handle is actively selected. Once selected, the handle can be moved freely while maintaining the pinch gesture. The design logic of the corner handles ensures that as one handle is adjusted, the handle on the directly opposite side of the Object Manipulation Box remains fixed, serving as an anchor. This mechanism allows for precise resizing of the objects while preserving their original positions within the scene.
Additionally, a uniform scaling feature is activated when the user turns their left hand sideways instead of up or down (Figure 5a and 5b). When this feature is enabled and the corner handle is selected, a line appears extending from the center of the Object Manipulation Box going through the selected corner handle. This line indicates the path along which the corner handle can move. The system calculates the nearest pinch location to this line to align the corner handle accurately, ensuring uniform scaling of the selected objects. While selecting multiple objects enables synchronized scaling, free scaling does not preserve the original aspect ratios of the objects (Figure 5c).
@@
When any of the elements for manipulating objects are held with the right hand with a pinch gesture, the orientation of the left hand dictates the toggle between free movement and snap-to-grid modes instead of single and multiple selection modes. Free movement is enabled when the left hand’s palm is facing upward, while snap-to-grid is activated when the palm is facing downward. The snap-to-grid functionality varies for movement, rotation, and scaling operations.
@@
Three photos labeled (a), (b), and (c) show a first-person view of a user looking at a wall. In the image (a) a solid cube and a hole sphere intersecting with the top corner of the cube is seen. The hole (or negative) sphere is seen with red color and with dotted transparent skin. In image (b) the left-hand wrist is visible while the cube and the sphere are selected and colored cyan. The user is about to poke a button named ”combine” with their right-hand index finger. In the image (c), the combination of the objects is visible where there is a round cut on the upper corner of the cube where the sphere used to intersect the cube.
Three photos labeled (a), (b), and (c) show the first-person view of the user looking at a wall with a vertical grid on the wall, holding their left-hand palm looking up, and manipulating a cube object. In image (a) user does a pinch motion to the movement handle on top of the cube. There are arrows appearing perpendicular to the sides of the cube with uniform steps. Similarly, in image (c), the same arrows are seen but for the scaling handle on the corner of the cube. In image (b), the user applies a pinch motion to the rotation handle on top of the cube and a circle with 15 degrees of steps parallel to the top side of the cube is seen.
Four images labeled (a), (b), (c), and (d) show a user’s first-person view using the system. In picture (a), user is using their left and right hand to pinch both ends of a digital line where in the middle of the line it says ”20.00 cm”. The user is measuring out the length of a game console (Nintendo Switch). In pictures (b), (c), and (d) user is manipulating a digital cube model. In the picture (b) user is applying a pinch motion to a small black handle on the middle of the edge of the cube. In picture (c), a numerical keyboard interface appears on the left hand in direction of the palm and it says ”10 cm” in the input field of the keyboard interface. In picture (d), the user has inputted ”20 cm” as the input and clicks the confirmation button. The cube’s width is doubled.

### v2

@@ 2.3. Novice Support Tools and Systems
Designing systems for novices requires consideration of simplicity, intuitiveness, and ease of learning. In the context of digital design and fabrication, this often translates into reducing the complexity of traditional CAD tools and making them more accessible. Tinkercad, for instance, has been successful in this regard by providing an entry-level CAD tool that embraces CSG modeling with simple geometric shapes and a drag-and-drop interface that is easy for beginners to grasp (Stemasov et al., 2021). Other systems for design and fabrication using simplified core building blocks include FlatFitFab (McCrae et al., 2014), which connects planar parts for a 3D model, and similarly kyub (Baudisch et al., 2019), which uses simpler block-based design. PARTs (Hofmann et al., 2018) addresses the challenges faced by non-experts in 3D modeling by introducing a framework for expressing and reusing design intent. CraftML (Yeh and Kim, 2018) uses declarative syntax where the code dictates the output structure. Lee et al. (Lee et al., 2018, 2016) use pose and movement detection for personalized furniture design, and Tian et al. simplify analog fabrication with woodcutter (Tian et al., 2018) and lathe (Tian et al., 2019) tools. Suzuki et al. (Suzuki et al., 2022) use virtual twins to visualize simulated behaviors when integrated with spatial references. Pfeuffer et al. (Pfeuffer et al., 2017) shows the use of virtual buttons on the wrists to access interactable menus.
@@
Generative design has also been instrumental in achieving desired 3D outcomes with minimal effort. Some tools allow users to design and fabricate 3D models based on 2D sketches (Saul et al., 2010; Li et al., 2020; Kazi et al., 2017; Johnson et al., 2012; Igarashi et al., 1999), thereby reducing the complexity associated with directly creating models in 3D space. Other ways of generative design utilize speech inputs (Ballagas et al., 2019) and constrained parametric design (Shugrina et al., 2015). Furthermore, retrieving existing designs (Stemasov et al., 2020; Liang et al., 2022) and remixing (Follmer and Ishii, 2012; Follmer et al., 2010; Stemasov et al., 2023b; Roumen et al., 2018) enhance accessibility while balancing expressivity (Stemasov et al., 2021). ShapeFindAR (Stemasov et al., 2022) enables searching for future physical artifacts through spatial queries and text.
Among novice-oriented tools, TinkerXR takes inspiration from Tinkercad’s simplicity and accessibility, especially its use of CSG modeling. By adopting an approachable modeling paradigm like CSG, combining it with spatial interaction features, and integrating it into an end-to-end design and fabrication workflow, TinkerXR aims to lower the entry barrier for beginners and introduce new opportunities for creating designs directly in their context.
@@ 3. DESIGN RATIONALE
When developing TinkerXR, several key design considerations were prioritized to ensure the system was both effective and intuitive for novice users. The main considerations included are user-friendly design and interaction, workspace awareness, and seamless integration:
@@ 3.1. User-Friendly Design and Interaction
To enhance the user experience in TinkerXR, we prioritize both simplicity in design and intuitive interaction. Our approach combines straightforward design logic with immersive user interactions to create an accessible and engaging environment for all users. The system’s design and interaction features are crafted to ensure that even novice users can seamlessly create and manipulate models, while advanced users benefit from an efficient design process.
@@
TinkerXR is designed with a focus on simplicity to ensure a novice-friendly user experience. By utilizing CSG operations, users can easily create complex models by combining basic shapes (e.g., cuboids and spheres) through simple boolean operations such as union and subtraction. This approach reduces complexity, allowing users to design without managing technical parameters. The system categorizes shapes into solid and hole types, enabling the addition or removal of material to create more advanced designs effortlessly.
@@
Interactivity in TinkerXR is driven by hand gestures and movement, recognized by Quest’s tracking system. These gestures allow users to manipulate objects directly, making the design process intuitive and engaging. Building upon the methods outlined in prior works (Kim et al., 2005; Weichel et al., 2014; Stemasov et al., 2024), TinkerXR lets users interact with the elements directly with their hands and without the need for external tools other than the headset. For instance, a pinching motion allows users to grab and move objects, while wrist rotation toggles features such as snap-to-grid, multi-selection, and uniform scaling. Moving the weak hand along a single global axis locks object manipulation to that axis, ensuring precision and control. Wrist-mounted menus provide an intuitive interface with buttons for quick navigation, including shortcuts for pinch gestures using fingers other than the index (e.g., select/deselect all). Visual and auditory feedback confirms actions, further enhancing the system’s usability and creating an immersive experience.
@@ 3.2. Workspace Awareness
TinkerXR’s real-world integration is achieved through sophisticated scene recognition and depth occlusion technologies. These allow the system to not only understand the environment but also intelligently interact with it.
Four photos labeled (a), (b), and two in (c). The first two photos show the first-person view of the user looking at a room with a table inside. The other two photos show three small decorative cacti on a table with a digital pyramid model behind them. In the image (a), three yellow rays come out from the user’s right hand, first hitting the floor where a text appearing above says ”select: floor”, second hitting the top of the table where a text appearing above says ”select: table”, and the third hitting the wall where a text appearing above says ”select: wall”. There is a white outline around the table. In image (b) the room is seen from the same angle as an image (a), but this time there are grids on top of the floor, table, and wall. In the first image of (c), the digital pyramid model partially blocks the view of the cacti even though it is placed behind them while in the second image of (c) the cacti are fully visible.
@@
It is essential for the system to recognize and adapt to its physical surroundings by using real-time environment mapping to adjust designs according to available space and nearby objects. The system can identify furniture and other objects, categorizing them as potential design spaces or obstacles, such as a desk, wall, or floor. This recognition ensures accurate placement and scaling of models within the environment. For instance, the system’s awareness of the selected workspace allows it to overlay a grid layout, enabling users to utilize snap-to-grid features and visualize object projections on the workspace as a reference during the design process. Additionally, the system supports the creation of reference objects, which are simple wireframe geometries representing physical items (e.g., a plant pot when designing a plant pot holder). These reference objects enable users to place the workspace grid plane on them, facilitating designs that seamlessly integrate with real-world items.
@@
It is also essential for users to understand their surroundings for context awareness while designing. While AR functionalities significantly aid in this area, models rendered in front of every world object can detract from the realistic nature we aim to achieve in the design task. By understanding the depth of objects in the environment, TinkerXR can occlude parts of a model that are behind physical objects, enhancing the realism and immersion of the design experience while elevating the user’s perception of the design as part of their real environment.
@@ 3.3. Seamless Integration
The transition between designing in AR and preparing a model for fabrication (in our case, 3D printing) should be fluid and straightforward, encouraging users to iterate on designs quickly and efficiently.
@@
In TinkerXR, we have consolidated design and fabrication functionalities into a single system, streamlining the entire workflow. Users can design their models and select printing options (e.g., layer height, infill pattern, support structure) directly within the AR environment, eliminating the need to transfer designs between separate tools. This integration simplifies the process by allowing users to handle all aspects of design and preparation within a unified interface, thereby reducing the complexity traditionally associated with using multiple software packages.
@@ 4. Implementation
TinkerXR was implemented for Meta Quest 3. The system code is written in C# with Unity Version 2022.3.22f1 LTS666https://unity.com/releases/2022-lts, Accessed: 19.07.2024. The system utilizes open-source resources provided by Meta to support various features. Specifically, we used Meta XR All-in-One SDK777https://developer.oculus.com/downloads/package/meta-xr-sdk-all-in-one-upm/, Accessed: 19.07.2024, which includes libraries for hand tracking, room and furniture detection (Scene API888https://developer.oculus.com/documentation/native/android/mobile-scene-api-ref/, Accessed: 2.08.2024), passthrough features, and other essential XR development tools for Unity. Additionally, we incorporated Meta’s Depth API999https://github.com/oculus-samples/Unity-DepthAPI, Accessed: 19.07.2024 library for the depth occlusion feature and Unity’s built-in audio engine for the spatial audio utilities. The subsequent sections will outline the envisioned procedural steps of using TinkerXR with its features, from (1) selecting the workspace to (2) designing artifacts, and finally to (3) fabrication of those artifacts.
@@ 4.1. Workspace Selection and Modification
TinkerXR system enables users to select their desired workspace, which can vary from the top of a desk to the floor, walls, door frames, and other flat surfaces.
@@ 4.1.2. Workspace Selection
After scanning, users can interact in TinkerXR and view the highlighted selectable workspaces. This is possible with the headset’s built-in room recognition feature and MRUK framework101010https://developer.oculus.com/documentation/unity/unity-mr-utility-kit-overview/, Accessed: 20.07.2024. Then, users use their dominant hand to direct a ray toward the surfaces of furniture or room boundaries, which displays the name of the planned workspace it intersects (Figure 2a). An index finger pinch gesture is then employed to confirm the selection.
@@ 4.2.1. Object Creation
Users can choose from seven types of primitive 3D geometries (cube, sphere, cylinder, capsule, triangular prism, pyramid, and cone) using buttons on their weak hand wrist (Figure 4a). The color of the button turns blue when the finger hovers close to the button. This gives the user feedback on which button is about to be pressed. Upon pressing a button, the selected object appears on the user’s weak hand (Figure 4b and 4c).
Five photos labeled (a), (b), (c), (d), (e) show the first-person view of a user in front of a wall interacting with their hand and the digital objects. In images (a), (b), and (c) user have extended their weak hand to their eye level and looking at the inside of their wrist. There are 12 white square buttons on the wrist in a three-by-four grid. Seven of the buttons show primitive geometric objects, namely cube, cylinder, sphere, capsule, triangular prism, pyramid, and cone. In the image (a), the user is hovering his dominant hand index finger on their wrist. In the image (b), the user pokes the cube button with their dominant hand index finger and the button turns green and a cube appears above the weak hand. In the image (c), the user does the same with a sphere next to the cube. In image (d), user does a pinch motion while their dominant hand hand is directed toward the created cube and their weak hand is oriented palm-down. The cube’s color is changed from blue to cyan and a white grid is appearing on the cube. Similarly, in the image (e), the user does the same pinch motion for the sphere, this time with the weak hand oriented palm-up. Both of the objects turn to cyan, with white grid appearing on the sphere instead of the cube.
@@ 4.2.2. Object Highlighting and Selection
During interactions, users may encounter multiple instances of objects, necessitating methods for selecting and manipulating specific items.
@@
A pinch gesture, i.e., thumb and index finger touching, selects a highlighted object (Figure 4d), turning it bright cyan and emitting a directional sound. A surrounding Object Manipulation Box appears, enabling movement, resizing, and rotation.
@@ 4.2.3. Single and Multiple Selection
Selection modes toggle based on weak-hand orientation: palm down for single selection (default, Figure 4d) and palm up for multiple selection (Figure 4e).
@@
In the single selection mode, only one object is active; selecting another deselects the current object. As this is considered the default mode, the palm-down orientation is chosen as it is the resting position of the hand.
@@
This mode enables group manipulation, i.e., the equivalent of pressing the control key on a keyboard in desktop interfaces. Selecting an object that is not already selected adds it to the list of selected objects. Conversely, selecting an already selected object removes it from the list. Multiple selected objects can be manipulated together at once.
@@ 4.2.4. Object Manipulation
The Object Manipulation Box facilitates object movement, resizing, and rotation.
Six photos labeled (a), (b), (c) each label corresponding to two photos show a cube being manipulated with their dominant hand from the first-person view of the user in front of a wall. The first pair of pictures labeled (a) shows the user pinching the movement handle on top of the cube in the first picture and dragging their dominant hand and the object to the right in the second picture. The second pair of pictures labeled (b) shows the user pinching the rotation handle next to the cube in the first picture and rotating their dominant hand wrist and the cube in the second picture. The third pair of pictures labeled (c) shows the user pinching the corner handle at the top right corner of the cube in the first picture and moving their hand to the right and the cube stretching to the right in the second picture.
@@
A movement handle, i.e., gizmo, above the objects follows the user’s pinch gesture, which allows for freehand movement (Figure 5a).
@@
Three axis-aligned rotation handles allow rotation around the x, y, and z axes. Grasping a handle with a pinch gesture rotates objects like a knob (Figure 5b).
@@
Corner handles at each vertex of the Object Manipulation Box enable resizing. Adjusting a handle moves it freely while the opposite handle remains fixed, which ensures precise control (Figure 5c).
Uniform scaling activates when the weak hand is sideways (Figure 6a, 6b). A guiding line extends from the box center through the selected corner, ensuring accurate scaling along this line. Multiple selections allow synchronized scaling, while free scaling does not preserve aspect ratios (Figure 6c).
Three photos labeled (a), (b), and (c) show a first-person view of a user looking at a wall and interacting with a complex 3D abstract design. In images (a) and (b), the user has their weak hand’s palm looking right and they apply a pinching motion to the corner handle on the upper right side of the manipulation box. A pink ray is seen coming from the middle and going through the selected corner handle. The user brings their right hand in the second picture closer to the center of the manipulation box, resulting in a smaller object with the same aspect ratio. Image (c) shows the user dragging the same corner handle with their weak hand palm looking down. The resulting object is compressed vertically.
@@
When any of the elements for manipulating objects are held with the dominant hand with a pinch gesture, the orientation of the weak hand dictates the toggle between free movement and snap-to-grid modes instead of single and multiple selection modes. Free movement is enabled when the weak hand’s palm faces upward, while snap-to-grid is activated when the palm faces downward. The functionality adapts to movement, rotation, and scaling operations.
Three photos labeled (a), (b), and (c) show the first-person view of the user looking at a wall with a vertical grid on the wall, holding their weak hand palm looking up, and manipulating a cube object. In image (a) user does a pinch motion to the movement handle on top of the cube. There are arrows appearing perpendicular to the sides of the cube with uniform steps. Similarly, in image (c), the same arrows are seen but for the scaling handle on the corner of the cube. In image (b), the user applies a pinch motion to the rotation handle on top of the cube and a circle with 15 degrees of steps parallel to the top side of the cube is seen.
@@
4.2.5. Axis Locking
The axis lock tool in TinkerXR allows users to easily constrain object manipulation to specific axes or planes, enhancing precision during design. Users can activate the axis lock tool by performing an index finger pinch gesture with their weak hand. Moving their hand in the desired global direction triggers a visual cue in the form of a colored edge (Figure 8a) or a side (if moved diagonally) of the axis lock tool, representing the active axis or plane of manipulation. This interaction ensures that users can limit movement or resizing action along a single axis or within a plane (Figure 8b and 8c) without needing to interact with additional buttons or menus. A single tap using the pinch gesture with their weak hand resets the axis lock, allowing users to resume unconstrained manipulation.